# How Many Ratings per Item are Necessary for Reliable Significance Testing?

### v2

@@ 1 Introduction
On the model side, response variance can come from stochastic inference, which is responsible for the creative power of foundation models, such as LLMs. It can also come from race conditions Shanmugavelu et al. (2024), mixtures of experts Shazeer et al. (2017)), Monte Carlo dropout Gal and Ghahramani (2016), and ensembling Lakshminarayanan et al. (2017).
@@ 2 Related Work
Wein et al. (2023) investigate p-value sensitivity of both metrics and test-set sampling methods in hypothesis testing, which therefore can affect the power analysis. While the latter did not turn out to be important in our study, metrics did. Clearly, different metrics (e.g., mean absolute error vs Spearman rank-correlation) will produce different scores for the same matrix of responses, so it stands to reason that any comparison will have different p-values for different metrics. They model a metric as a function $\Gamma(M,G)$, where $M$ is a matrix of model predictions which returns a score for $M$. We assume $\Gamma$ is given here but focus on the best performing of these metrics in experiments. Homan et al. (2024) initiates a study of the trade-off between number of items and responses using a toy simulator. By contrast, we use real datasets to investigate these trade-offs and perform experiments that shed light on the mechanism for how response variance provides statistical significance.
@@ 4 Methods
Our main contribution is a human-in-the-loop process that allows one to (1) estimate the amount of data in terms of items, $N$, and responses per item, $K$, needed to detect, with high confidence, a difference of performance according to metric $\Gamma$ of at least $\epsilon$; and (2) compute p-values for existing experimental data comparing the performance of two models against gold data. Note that when the amount of experimental data is insufficient we can fit the data to a parameterized model and perform (1) to rerun the experiments with a sufficiently large dataset. It is precisely this use case that our experiments address.
@@
We then use a simulator to generate new gold responses the the same (fitted) distribution as $G$. We us same given distribution to generate data for both $A$ and $G$, so that $A$ represents an ideal model for $G$. We add perturbation (governed by $\epsilon$) to this distribution to generate data for $B$. This ensures that model $A$ performs better than model $B$ with respect to $G$ under almost any metric, and that “ground-truth” $p$-values should converge to zero as $\epsilon$, $N$, and/or $K$ increase. The simulator then estimates $p$-values (or, in the case of power analysis $1-\beta$) based on a large number of repetitions $b$. Typically $b=10000$, although power analysis requires two levels of repetitions: one to generate a distribution over effect sizes and one to estimate the p-value, given the effect size. We report the number of repetitions in the figures associated with each of our results (Figure 3).
@@ 5.1 Data
Note that we chose the folded and truncated normal and triangular distributions for these datasets based on visually matching histograms of the responses of each dataset, as described in Section 4. We can use any family of distributions we like, ı.e., they need not be any flavor of normal distribution, as long as there are algorithmically feasible ways of fitting them to the data.
Figure 1 illustrates goodness-of-fit for simulations of two datasets used in this paper. Details and results for additional datasets can be found in Appendix C.
@@ 5.2 Results
(a) Stanford Toxicity
(b) MultiDomain Agreement
(a) Toxicity ($\epsilon=0.05$): MultiStage vs “flat” bootstrap
(b) MultiDomain ($\epsilon=0.1$): MultiStage vs “flat” bootstrap
@@
Upon publication, we will release the code used to run our experiments to the public. We used the Python libraries NumPy, Pandas, and SciPy, versions 2.2.3, 2.2.1, and 1.13.1, respectively. Our experiments took various times to run, with the longest experiments (producing any of the points in our figures) running approximately nine hours.
@@ 7 Conclusion
Appendix A Fitting the Simulator to Real Data
However, because we do have enough responses over all items, we do choose for the first stage specific distributions for each dataset that, paired with the second stage described above, fit the data.
Figure 6: Shapes possible with censored normal.
We used the censored normal distribution for $\mathcal{N}$, which assumes a latent continuous distribution that is not observed exactly but measured to within intervals, including left and right intervals which pool (not truncate) the smallest and largest values, respectively. This provides support for head and/or tail bias; Figure 6 illustrates a variety shapes that this distribution can capture. For example, items in the Stanford Toxicity dataset (see Section 5.1) rated at either extreme (either “not toxic” or “extremely toxic”) tend to have more agreement among raters. We use distributions fitted to each dataset from distribution families tailored to each dataset. This involves visualizing the distributions of response means and standard deviations of the item responses in each dataset to get a sense of what they look like and then choosing a parameterized family of distributions to fit the data to. Figure 1 illustrates goodness-of-fit for simulations of two datasets used in this paper.
Appendix B Using the simulator to estimate $p$-values
B.1 Simulator
@@
The simulator takes input parameters $N$ and $K$, along with perturbation parameter $\epsilon$. In the first stage, it randomly chooses hyperparameters $\mathbf{\theta}_{1},\ldots,\mathbf{\theta}_{N}\sim P_{items}$, each corresponding to an item $\mathbf{\theta}_{i}$, from a fixed distribution that serve as model parameters for the second stage. In the second stage, for each item $i$ we sample $K$ responses from a second distribution $P_{responses}(\mathbf{\theta}_{i})$. We do this for each of the datasets, respectively representing responses from gold annotations, $G^{N\times K}$, and two models, $A^{N\times K}$ and $B^{N\times K}$. The specific distributions that were used in our experiments were modeled from real datasets; for details see Appendix A.
@@
B.2 Estimating $p$-values
Given $N$, $K$, and $\epsilon$, p-values are estimated by drawing $b$ (bootstrap) resamples $S_{alt}=\langle G^{N\times K}_{1},A^{N\times K}_{1},B^{N\times K}_{1,\epsilon}\rangle,$ $\ldots,$ $\langle G^{N\times K}_{b},A^{N\times K}_{b},B^{N\times K}_{b,\epsilon}\rangle$ for the alternative hypothesis according to the process described in Section B.1. Since the null hypothesis makes the assumption that the distributions of $A$ and $B$ are the same with respect to $G$, we construct $S_{null}$ by pooling the items from $A^{N\times K}$ and $B^{N\times K}$ and then independently sampling from this pool. When sampling responses from $A$, for each item $i$, we sample each response by sampling from $P_{responses}(\mathbf{\theta}_{i})$, where $\mathbf{\theta}_{i}=(\mu_{i},\sigma_{i})$. Sampling responses from $B$ is similar but we first choose $\delta_{i}\sim Unif(-\epsilon,\epsilon)$ and then sample from $P_{responses}(\mathbf{\theta}_{i})$, where $\mathbf{\theta}_{i}=(\mu_{i}+\delta_{i},\sigma_{i})$.
@@ B.2 Estimating $p$-values
Appendix C Results on Additional Datasets
@@
Dataset Stat $\Gamma_{\rm MAE}$ $\Gamma_{\rm Wins}$ $\Gamma_{\rm MEMD}$
NK 10000 25000 25000
Toxicity p-value 0.047 0.026 0.011
K 100 100 100
$\Delta$ 0.007 0.227 0.546
NK 100000 - 100000
MultiDomain p-value 0.019 - 0.033
K 100 - 100
$\Delta$ 0.006 - 0.195
NK 40000 100000 -
Amazon p-value 0.043 0.031 -
K 100 100 -
$\Delta$ 0.005 0.118 -
NK - - -
HS-Brexit p-value - - -
K - - -
$\Delta$ - - -
NK 40000 100000 40000
ConvAbuse p-value 0.020 0.037 0.033
K 100 100 100
$\Delta$ 0.009 0.093 0.279
NK 100000 - 100000
ArMIS p-value 0.025 - 0.047
K 100 - 100
$\Delta$ 0.005 - 0.169
NK 20000 40000 20000
MHS p-value 0.041 0.044 0.037
K 100 100 100
$\Delta$ 0.001 0.107 0.246

### v3

@@ 1 Introduction
On the model side, response variance can come from stochastic inference, which is responsible for the creative power of foundation models, such as LLMs. It can also come from race conditions Shanmugavelu et al. (2024), mixtures of experts Shazeer et al. (2017), Monte Carlo dropout Gal and Ghahramani (2016), and ensembling Lakshminarayanan et al. (2017).
@@ 2 Related Work
Wein et al. (2023) investigate p-value sensitivity of both metrics and test-set sampling methods in hypothesis testing, which therefore can affect the power analysis. While the latter did not turn out to be important in our study, metrics did. Clearly, different metrics (e.g., mean absolute error vs Spearman rank-correlation) will produce different scores for the same matrix of responses, so it stands to reason that any comparison will have different p-values for different metrics. They model a metric as a function $\Gamma(M,G)$, where $M$ is a matrix of model predictions which returns a score for $M$. We assume $\Gamma$ is given here but focus on the best performing of these metrics in experiments. Homan et al. (2024) initiates a study of the trade-off between number of items and responses using a toy simulator. By contrast, we use real datasets to investigate these trade-offs and perform experiments that shed light on the mechanism for how response variance provides statistical significance. Recently, in a follow-up paper Pandita et al. (2025) to this work, we modeled categorical datasets using a Bayesian approach and examined the optimization problem of allocating a fixed human annotation budget ($N\times K$). The current work provides the foundational evidence, for regression models, that increasing responses per item ($K$) is often more critical for significance than increasing the number of items ($N$).
@@ 4 Methods
Our main contribution is a human-in-the-loop process that allows one to (1) estimate the amount of data in terms of items, $N$, and responses per item, $K$, needed to detect, with high confidence, a difference of performance according to metric $\Gamma$ of at least $\epsilon$; and (2) compute p-values for existing experimental data comparing the performance of two models against gold data. Note that when the amount of experimental data is insufficient, we can fit the data to a parameterized model and perform (1) to rerun the experiments with a sufficiently large dataset. It is precisely this use case that our experiments address.
@@
We then use a simulator to generate new gold responses as the same (fitted) distribution as $G$. We use the same given distribution to generate data for both $A$ and $G$, so that $A$ represents an ideal model for $G$. We add perturbation (governed by $\epsilon$) to this distribution to generate data for $B$. This ensures that model $A$ performs better than model $B$ with respect to $G$ under almost any metric, and that “ground-truth” $p$-values should converge to zero as $\epsilon$, $N$, and/or $K$ increase. The simulator then estimates $p$-values (or, in the case of power analysis $1-\beta$) based on a large number of repetitions $b$. Typically $b=10000$, although power analysis requires two levels of repetitions: one to generate a distribution over effect sizes and one to estimate the p-value, given the effect size. We report the number of repetitions in the figures associated with each of our results (Figure 4).
@@ 5.1 Data
5.2 Fitting the Simulator to Real Data
Figure 1: Shapes possible with censored normal.
However, because we do have enough responses over all items, we do choose for the first stage specific distributions for each dataset that, paired with the second stage described above, fit the data. We used the censored normal distribution for $\mathcal{N}$, which assumes a latent continuous distribution that is not observed exactly but measured to within intervals, including left and right intervals which pool (not truncate) the smallest and largest values, respectively. This provides support for head and/or tail bias; Figure 1 illustrates a variety shapes that this distribution can capture. For example, items in the Stanford Toxicity dataset (see Section 5.1) rated at either extreme (either “not toxic” or “extremely toxic”) tend to have more agreement among raters.
We use distributions fitted to each dataset from distribution families tailored to each dataset. Note that we chose the folded and truncated normal and triangular distributions for these datasets based on visually matching histograms of the responses of each dataset, as described in Section 4. We can use any family of distributions we like, i.e., they need not be any flavor of normal distribution, as long as there are algorithmically feasible ways of fitting them to the data. Figure 2 illustrates goodness-of-fit for simulations of two datasets used in this paper. Details about computing p-values and results for additional datasets can be found in Appendix A and B.
(a) Stanford Toxicity
(b) MultiDomain Agreement
@@ 5.3 Results
(a) Toxicity ($\epsilon=0.05$): MultiStage (left) vs “flat” bootstrap (right)
(b) MultiDomain ($\epsilon=0.1$): MultiStage (left) vs “flat” bootstrap (right)
@@
We utilize the Variance Estimation Toolkit (VET) [^1] https://github.com/google-research/vet to run our experiments. We used the Python libraries NumPy, Pandas, and SciPy, versions 2.2.3, 2.2.1, and 1.13.1, respectively. Our experiments took various times to run, with the longest experiments (producing any of the points in our figures) running approximately nine hours.
@@ 7 Conclusion
Appendix A Using the simulator to estimate $p$-values
A.1 Simulator
@@
The simulator takes input parameters $N$ and $K$, along with perturbation parameter $\epsilon$. In the first stage, it randomly chooses hyperparameters $\mathbf{\theta}_{1},\ldots,\mathbf{\theta}_{N}\sim P_{items}$, each corresponding to an item $\mathbf{\theta}_{i}$, from a fixed distribution that serve as model parameters for the second stage. In the second stage, for each item $i$ we sample $K$ responses from a second distribution $P_{responses}(\mathbf{\theta}_{i})$. We do this for each of the datasets, respectively representing responses from gold annotations, $G^{N\times K}$, and two models, $A^{N\times K}$ and $B^{N\times K}$. The specific distributions that were used in our experiments were modeled from real datasets; for details see Appendix 5.2.
@@
A.2 Estimating $p$-values
Given $N$, $K$, and $\epsilon$, p-values are estimated by drawing $b$ (bootstrap) resamples $S_{alt}=\langle G^{N\times K}_{1},A^{N\times K}_{1},B^{N\times K}_{1,\epsilon}\rangle,$ $\ldots,$ $\langle G^{N\times K}_{b},A^{N\times K}_{b},B^{N\times K}_{b,\epsilon}\rangle$ for the alternative hypothesis according to the process described in Section A.1. Since the null hypothesis makes the assumption that the distributions of $A$ and $B$ are the same with respect to $G$, we construct $S_{null}$ by pooling the items from $A^{N\times K}$ and $B^{N\times K}$ and then independently sampling from this pool. When sampling responses from $A$, for each item $i$, we sample each response by sampling from $P_{responses}(\mathbf{\theta}_{i})$, where $\mathbf{\theta}_{i}=(\mu_{i},\sigma_{i})$. Sampling responses from $B$ is similar but we first choose $\delta_{i}\sim Unif(-\epsilon,\epsilon)$ and then sample from $P_{responses}(\mathbf{\theta}_{i})$, where $\mathbf{\theta}_{i}=(\mu_{i}+\delta_{i},\sigma_{i})$.
@@ A.2 Estimating $p$-values
Appendix B Results on Additional Datasets
@@
Dataset Stat $\Gamma_{\rm MAE}$ $\Gamma_{\rm Wins}$ $\Gamma_{\rm MEMD}$
NK 10000 25000 25000
Toxicity p-value 0.047 0.026 0.011
K 100 100 100
$\Delta$ 0.007 0.227 0.546
NK 100000 - 100000
MultiDomain p-value 0.019 - 0.033
K 100 - 100
$\Delta$ 0.006 - 0.195
NK 40000 100000 -
Amazon p-value 0.043 0.031 -
K 100 100 -
$\Delta$ 0.005 0.118 -
NK - - -
HS-Brexit p-value - - -
K - - -
$\Delta$ - - -
NK 40000 100000 40000
ConvAbuse p-value 0.020 0.037 0.033
K 100 100 100
$\Delta$ 0.009 0.093 0.279
NK 100000 - 100000
ArMIS p-value 0.025 - 0.047
K 100 - 100
$\Delta$ 0.005 - 0.169
NK 20000 40000 20000
MHS p-value 0.041 0.044 0.037
K 100 100 100
$\Delta$ 0.001 0.107 0.246